OpenStudy (anonymous):

A rectangle has an area of 32 square feet and a perimeter of 24 feet. What is the shortest of the side lengths, in feet, of the rectangle? F. 1 G. 2 H. 3 J. 4 K. 8

OpenStudy (anonymous):

I got J. Is there a trick or shortcut to solving problems like this ?

OpenStudy (cggurumanjunath):

A = L * B = 32 P = 2 ( L + B ) = 24 LB=32 B=32/L 2(L+B)=24 2(L+(32/L))=24 L+((32/L)) =12 L^2+32=12L

OpenStudy (cggurumanjunath):

L^2-12L+32=0 L^2-4L-8L+32=0 L(L-4) -8(L-4) =0 (L-4)(L-8)=0 L=4;L=8 L=4 IS THE SHORTEST !

OpenStudy (cggurumanjunath):

SO U R RIGHT @katrinamarie0803 ! J=4 IS THE ANSWER .